OpenStudy (appleduardo):

whats the integral for x^4 / (1-x) ??

terenzreignz (terenzreignz):

\[\huge \int \frac{x^4}{1-x}dx\]

OpenStudy (appleduardo):

yep, but how can I solve it?

terenzreignz (terenzreignz):

It's actually quite easy, but incredibly tedious. Use u-substitution. When in doubt, attempt to let u = the denominator of a rational expression... chances are, that's the one...

OpenStudy (appleduardo):

what u mean with rational expression?

terenzreignz (terenzreignz):

Fraction. Fancy word for fraction.

OpenStudy (appleduardo):

so u say that Ive to use u=x^4, du/dx=4x^3 so dx=du/4x^3 then: \[\frac{ u }{ 1-x }\frac{ du }{ 4x^3 }\]

OpenStudy (appleduardo):

??

terenzreignz (terenzreignz):

Unfortunately not that simple. u was in your denominator, wasn't it?

OpenStudy (appleduardo):

oh yeep ure right, so its: u=(1-x) du/dx= -1 dx= du/-1 and then: \[\int\limits_{}^{}\frac{ x^4 }{ u}*\frac{ du }{ -1 } = -\int\limits_{}^{}\frac{ x^4 }{ u}*du\]

OpenStudy (appleduardo):

??

terenzreignz (terenzreignz):

Okay, much better. But you cannot solve this integral without expressing \(x^4\) in terms of u.

OpenStudy (appleduardo):

so.. what can I do??

terenzreignz (terenzreignz):

Well \[\large u = 1-x\] \[\large x = 1-u\] \[\huge x^4 = (1-u)^4\]

OpenStudy (appleduardo):

wow! so uhm,, what do I have to do now?

terenzreignz (terenzreignz):

Expand.

terenzreignz (terenzreignz):

\[\large (1-u)^4 = u^4 -4u^3 +6u^2 -4u +1\]

OpenStudy (appleduardo):

so now I have: \[\frac{ x^4 }{ u^4 - 4u^3 + 6u^2 - 4u + 1 }\] but at this point is u still = to (1-x) ? sorry if this sounds silly or so, but I got a little confused when u got (1-u)^4

terenzreignz (terenzreignz):

No... remember, you started with \[\huge \int \frac{x^4}{1-x}dx\]And you let u = 1-x, work from there, and substitute.

OpenStudy (appleduardo):

mm so what I = to "u" then? :/

terenzreignz (terenzreignz):

Shun being spoonfed, @appleduardo ... :P \[\large u = 1-x\]\[\large du = -dx\]\[\large dx = -du\]\[\large x = 1-u\]\[\large x^4=(1-u)^4\]